# Towards a robust frequency-domain analysis: Spectral Rényi divergence revisited

### v1

@@ Abstract
This paper studies a specific category of statistical divergences for spectral densities of time series: the spectral $\alpha$-Rényi divergences, which includes the Itakura–Saito divergence as a subset. While the spectral Rényi divergence has been acknowledged in past works, its statistical attributes have not been thoroughly investigated. The aim of this paper is to highlight these properties. We reveal a variational representation of spectral Rényi divergence, from which the minimum spectral Rényi divergence estimator is shown to be robust against outliers in the frequency domain, unlike the minimum Itakura–Saito divergence estimator, and thus it delivers more stable estimate, reducing the need for intricate pre-processing.
@@
Frequency domain estimation; Robust inference; Spectral density; Statistical divergence
@@ 1 Introduction
Frequency-domain analysis of time series data has been conducted in many applied fields. Central to this analysis is the (power) spectral density, a crucial element whose estimation has garnered significant attention. For instance, in seismology, earthquake source parameters are delineated through the spectral parameter estimation [3, 20]. In audio signal processing, various methodologies have been developed via spectral density estimation to achieve signal separation [10].
@@
In this paper, we take another look at a particular class of spectral divergence: the spectral $\alpha$-Rényi divergences [17, 4, 5]. Notably, the spectral Rényi divergence encompasses the Itakura–Saito divergence as a special instance. Though this class has been previously mentioned in the extended discussion of ”Pinsker’s information theoretic justification of the Itakura-Saito distortion measure” [13] in the existing literature, its statistical characteristics have been largely overlooked.
The primary aim of this paper is to elucidate the statistical properties of this class. Specifically, we find a variational representation of spectral Rényi divergence and demonstrate that, from this representation, the minimum spectral Rényi divergence estimator possesses robustness against outliers in the frequency domain, a quality that the minimum Itakura–Saito divergence estimator does not have. These outliers in the frequency domain often emerge due to insufficient detrending [6, 7, 12]. Yet, when using the Itakura–Saito divergence, practitioners should exercise greater caution regarding detrending and other pre-processing steps. In comparison, spectral Rényi divergences offer more stable estimation results without necessitating specific meticulous pre-processing.
@@ 2 Spectral Rényi divergence
This section introduces a class of spectral Rényi divergences and its properties. Let $\mathcal{S}:=\{S(\cdot):\int_{-\pi}^{\pi}S(\omega)\,\mathrm{d}\omega<\infty,\,S(\omega)>0\,\text{for all $\omega\in(0,\pi)$},S(\omega)=S(-\omega)\}$. Fix $S\in\mathcal{S}$ and $\widetilde{S}\in\mathcal{S}$ arbitrarily. Let $p_{n}(\mathbf{x}_{n})$ and $\widetilde{p}_{n}(\mathbf{x}_{n})$, $\mathbf{x}_{n}\in\mathbb{R}^{n}$, be the $n$-th order probability densities of two zero-mean stationary Gaussian processes with spectral densities $S$ and $\widetilde{S}$, respectively. We assume the sampling rate is equal to 1.
@@
$D_{\alpha}[\,S\,:\,\widetilde{S}\,]:=\frac{1}{2\pi(1-\alpha)}\int_{-\pi}^{\pi}[\log\{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)\}-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)]\,\mathrm{d}\omega.$ (1)
@@
$D_{1}[\,S\,:\,\widetilde{S}\,]:=\lim_{\alpha\to 1^{-}}D_{\alpha}[\,S\,:\,\widetilde{S}\,]=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left\{\left(\frac{\widetilde{S}(\omega)}{S(\omega)}\right)^{-1}-1+\log\left(\frac{\widetilde{S}(\omega)}{S(\omega)}\right)\right\}\,\mathrm{d}\omega,$ (2)
@@
The name of this class comes from the fact that the class is induced by the limit of the probabilistic Rényi divergence between stationary Gaussian processes, which was pointed out by [17].
Theorem 1 (Proposition 8.29 of [17]).
For $\alpha\in(0,1]$, the following holds:
@@
$D_{\alpha}[\,p_{n}\,:\,\widetilde{p}_{n}\,]:=\frac{1}{\alpha-1}\log\int_{\mathbb{R}^{n}}\left(\frac{p_{n}(\mathbf{x}_{n})}{\widetilde{p}_{n}(\mathbf{x}_{n})}\right)^{\alpha-1}p_{n}(\mathbf{x}_{n})\,\mathrm{d}\mathbf{x}_{n}$
@@
$D_{\alpha}^{(n)}[\,S\,:\,\widetilde{S}\,]=\frac{1}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left[\log\left\{\alpha\widetilde{S}\left(\omega\right)+(1-\alpha)S\left(\omega\right)\right\}-\alpha\log\widetilde{S}\left(\omega\right)-(1-\alpha)\log S\left(\omega\right)\right],$
where $\Omega_{n}:=\{2\pi(t/n):t=-\lceil n/1\rceil+1,\ldots,-1,0,1,\ldots,\lfloor n/2\rfloor\}$. By the convergence of the Riemann sums, $D^{(n)}_{\alpha}$ converges to $D_{\alpha}$ as $n$ goes to infinity.
2.1 Variational representation of spectral Rényi divergences
Here we introduce the variational representation of spectral Rényi divergences that have not been investigated in the literature:
@@ 2.1 Variational representation of spectral Rényi divergences
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left\{\frac{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)}{S^{\prime}(\omega)}-1+\log S^{\prime}(\omega)-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)\right\}\,\mathrm{d}\omega$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left\{\frac{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)}{S^{\prime}(\omega)}-1+\log\frac{S^{\prime}(\omega)}{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)}\right\}\,\mathrm{d}\omega$
$\qquad+\frac{1}{2\pi}\int_{-\pi}^{\pi}[\log\{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)\}-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)]\,\mathrm{d}\omega$
@@
$\qquad+\frac{1}{2\pi}\int_{-\pi}^{\pi}[\log\{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)\}-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)]\,\mathrm{d}\omega.$
@@
3 Robustness against outliers in the frequency domain
A tantalizing feature of the spectral $\alpha$-Rényi divergence for $\alpha\in(0,1)$ is that it is robust with respect to outliers in the frequency domain, which has been overlooked in the literature. We investigate this property of the spectral Rényi divergence. Let $\mathbf{x}_{n}=(x_{1},\ldots,x_{n})^{\top}$ be the $n$-th order data from a possibly non-Gaussian time series $\{X_{t}\}_{t\in\mathbb{Z}}$ with a spectral density and let $\mathcal{S}_{\Theta}:=\{S_{\theta}\in\mathcal{S}:\theta\in\Theta\}$ be a parametric spectral model with a parameter space $\Theta\subset\mathbb{R}^{d}$. Consider the minimum spectral $\alpha$-Rényi divergence estimator $\widehat{\theta}_{\alpha}$ given as
@@ 3 Robustness against outliers in the frequency domain
Consider the following contaminated one $\widetilde{I}_{n}^{z,\omega^{*}}$ at discrete frequencies $\Omega_{n}$: for $\omega^{*}\in\Omega_{n}$ and $z>0$,
@@
This type of contamination is motivated by the case when periodic components are not suitably subtracted and are contaminated in observed time series [6, 7, 12], say, $x_{t}^{o}=x_{t}+\sqrt{8\pi z/n}\cos(At)$, $t=1,\ldots,n$ with original zero-mean time series $\{x_{t}\}_{t=1,2,\cdots,n}$. In this case, the periodogram of $x_{t}^{o}$ has approximately the form of
@@
where $\delta_{A}(\omega)=1$ if $\omega=A$ and 0 otherwise. In what follows, we show that (1) the minimum Itakura–Saito divergence estimator is sensitive to outliers in the frequency domain, and (2) the minimum spectral Rényi divergence estimator is robust against to it.
3.1 Robustness of the trajectory of the optimization
We will examine the effects of contamination in the frequency domain on both $\widehat{\theta}_{1}$ and $\widehat{\theta}_{\alpha}$.

### v3

# On robustness of Spectral Rényi divergence

@@ Abstract
This paper studies a specific category of statistical divergences for spectral densities of time series: the spectral $\alpha$-Rényi divergences, which includes the Itakura–Saito divergence as a subset. The aim of this paper is to highlight both information-theoretic and statistical properties of spectral $\alpha$-Rényi divergences. We reveal the connection between the spectral $\alpha$-Rényi divergence and the $\gamma$-divergence in robust statistics, and a variational representation of spectral $\alpha$-Rényi divergence. Inspired by these results suggesting “robustness” of spectral $\alpha$-Rényi divergence, we show that the minimum spectral Rényi divergence estimate has a stable optimization path with respect to outliers in the frequency domain, unlike the minimum Itakura–Saito divergence estimator, and thus it delivers more stable estimate, reducing the need for intricate pre-processing.
@@
Frequency domain analysis; Optimization theory; Robust statistics; Spectral density; Statistical divergence; Time series analysis
@@ 1 Introduction
Frequency-domain analysis of time series data has been conducted in many applied fields. Central to this analysis is the (power) spectral density, a crucial element whose estimation has garnered significant attention. For instance, in seismology, earthquake source parameters are delineated through the spectral parameter estimation [5, 33]. In geodesy, temporal correlation of noises in the Global Navigation Satellite System is captured by spectral parameter estimation [17]. In audio signal processing, various methodologies have been developed via spectral density estimation to achieve signal separation [20].
@@
In this paper, we take another look at a particular class of spectral divergence: the spectral $\alpha$-Rényi divergences [29, 34, 32, 8, 9]. Notably, the spectral Rényi divergence encompasses the Itakura–Saito divergence as a special instance. In the probabilistic and information-theoretic literature, this class has been mentioned in the extended discussion of “Pinsker’s information theoretic justification of the Itakura-Saito distortion measure” [24]. In the statistical literature, the robustness of the spectral Rényi divergence against spectral peaks is pointed out [34], which has been utilized in clustering analysis of time series [32, 11].
The primary aim of this paper is to elucidate new properties of this class that shed light on the robustness in the spectral parameter estimation. Specifically, we show that (i) the $\gamma$-divergence [7] between probability densities of the processes leads to the spectral Rényi divergence, and (ii) the spectral Rényi divergence has a variational representation. Inspired by these new results, we explore (iii) an additional robustness property of the minimum spectral Rényi divergence estimate. The minimum spectral Rényi divergence estimate has been shown to be robust against time series outliers in the frequency domain [34, 32, 11]. Expanding on these findings, we show that the optimization path in obtaining the minimum spectral Rényi divergence estimate is stable in the presence of time series outliers, a quality that the minimum Itakura–Saito divergence estimator does not have.
Let us explain the implications of our results. The first result bridges spectral and probabilistic divergences, providing an information-theoretic justification of the spectral Rényi divergence. The second result gives a decomposition of spectral Rényi divergence in terms of the Itakura–Saito divergence. The first two results contribute to understanding the robustness of the spectral Rényi divergence. The third result has both theoretical and practical implications. Theoretically, it gives a novel connection between robust spectral analysis (c.f., [19]) and optimization theory. Developing robust statistics from an algorithmic or computational perspective has been a recent topic of discussion (c.f., [6]). Our results provide an additional perspective on robust statistics from a computational viewpoint, focusing on the stability of optimization paths. Practically, outliers in the frequency domain often emerge due to insufficient detrending [10, 12, 22]. Yet, when using the Itakura–Saito divergence, practitioners should exercise greater caution regarding detrending and other pre-processing steps. In comparison, spectral Rényi divergences offer more stable estimation results without necessitating specific meticulous pre-processing.
The structure of this paper is as follows. Section 2 delivers new formulae related to the spectral Rényi divergence (Theorem 1 and Theorem 2). Section 3 discusses the stability of optimization paths of the spectral Rényi divergence minimization (Theorem 3). Section 4 presents thorough numerical studies. Proofs of the results are presented in Appendices.
@@ 2 Spectral Rényi divergence
This section introduces a class of spectral Rényi divergences and its properties. Let $\mathcal{S}:=\{S(\cdot):\int_{-\pi}^{\pi}S(\omega)\,\mathrm{d}\omega<\infty,\,S(\omega)>0\,\text{for all $\omega\in(0,\pi)$},S(\omega)=S(-\omega)\}$. Fix $S\in\mathcal{S}$ and $\widetilde{S}\in\mathcal{S}$ arbitrarily. Let $p_{n}(\mathbf{x}_{n})$ and $\widetilde{p}_{n}(\mathbf{x}_{n})$, $\mathbf{x}_{n}\in\mathbb{R}^{n}$, be the $n$-th order probability densities of two zero-mean stationary possibly non-Gaussian processes with spectral densities $S$ and $\widetilde{S}$, respectively. We assume the sampling rate is equal to 1.
@@
$D_{\alpha}[\,S\,:\,\widetilde{S}\,]:=\frac{1}{2\pi(1-\alpha)}\int_{-\pi}^{\pi}[\log\{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)\}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)]\,\mathrm{d}\omega.$ (1)
@@
$D_{1}[\,S\,:\,\widetilde{S}\,]$ $:=\lim_{\alpha\to 1^{-}}D_{\alpha}[\,S\,:\,\widetilde{S}\,]$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left\{\left(\frac{\widetilde{S}(\omega)}{S(\omega)}\right)^{-1}-1+\log\left(\frac{\widetilde{S}(\omega)}{S(\omega)}\right)\right\}\,\mathrm{d}\omega,$ (2)
@@
The name of this class comes from the fact that the class is induced by the limit of the probabilistic Rényi divergence between stationary Gaussian processes, which was pointed out by [29]. That is, if two densities $p_{n}$ and $\widetilde{p}_{n}$ are Gaussian, the following holds for $\alpha\in(0,1]$:
@@
$D_{\alpha}[\,p_{n}\,:\,\widetilde{p}_{n}\,]:=\frac{1}{\alpha-1}\log\int_{\mathbb{R}^{n}}\left(\frac{p_{n}(\mathbf{x}_{n})}{\widetilde{p}_{n}(\mathbf{x}_{n})}\right)^{\alpha-1}p_{n}(\mathbf{x}_{n})\,\mathrm{d}\mathbf{x}_{n}.$
@@
$D_{\alpha}^{(n)}[\,S\,:\,\widetilde{S}\,]=\frac{1}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\Big{[}\log\left\{\alpha\widetilde{S}\left(\omega\right)+(1-\alpha)S\left(\omega\right)\right\}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\alpha\log\widetilde{S}\left(\omega\right)-(1-\alpha)\log S\left(\omega\right)\Big{]},$
where $\Omega_{n}:=\{2\pi(t/n):t=-\lceil n/2\rceil+1,\ldots,-1,0,1,\ldots,\lfloor n/2\rfloor\}$. By the convergence of the Riemann sums, $D^{(n)}_{\alpha}$ converges to $D_{\alpha}$ as $n$ goes to infinity.
2.1 Properties of spectral Rényi divergences
Here we present two new formulae that have not been investigated in the literature: (i) the asymptotic equivalence between the $\gamma$-divergence [7] and the spectral Rényi divergence; and (ii) the variational representation of spectral Rényi divergences.
We first state the asymptotic equivalence between the $\gamma$-divergence and the spectral Rényi divergence. Let $G_{\gamma}[\,p\,:\,q\,]$ be the $\gamma$-divergence between two probability densities $p$ and $q$:
$G_{\gamma}[\,p\,:\,q\,]=\frac{1}{\gamma(1+\gamma)}\log\int(p(x))^{1+\gamma}\,\mathrm{d}x-\frac{1}{\gamma}\log\int p(x)(q(x))^{\gamma}\,\mathrm{d}x$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{1+\gamma}\log\int(q(x))^{1+\gamma}\,\mathrm{d}x.$
Theorem 1 (The $\gamma$-divergence leads to spectral Rényi divergences).
Let $\alpha\in(0,1)$ and $S,\widetilde{S}\in\mathcal{S}$. If two densities $p_{n}$ and $\widetilde{p}_{n}$ are the $n$-th order probability densities of stationary Gaussian processes with spectral densities $S$ and $\widetilde{S}$, we have
$\lim_{n\to\infty}\frac{2}{n}G_{\gamma}[\,p_{n}\,:\,\widetilde{p}_{n}\,]=D_{\alpha}[\,S\,:\,\widetilde{S}\,],$
where $\gamma=\alpha^{-1}-1>0$.
The proof is given in Appendix A. It employs the direct evaluation of the $\gamma$-divergence and Szegö’s limit theorem[27, 15].
We next state the variational representation of the spectral Rényi divergence.
@@ 2.1 Properties of spectral Rényi divergences
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\Bigg{\{}\frac{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)}{S^{\prime}(\omega)}-1+\log S^{\prime}(\omega)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)\Bigg{\}}\,\mathrm{d}\omega$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\Bigg{\{}\frac{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)}{S^{\prime}(\omega)}-1+\log\frac{S^{\prime}(\omega)}{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)}\Bigg{\}}\,\mathrm{d}\omega$
$\,\,+\frac{1}{2\pi}\int_{-\pi}^{\pi}[\log\{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)\}-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)]\,\mathrm{d}\omega$
@@
$\,\,+\frac{1}{2\pi}\int_{-\pi}^{\pi}[\log\{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)\}-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)]\,\mathrm{d}\omega.$
@@
2.2 Robustness of the spectral Renyi divergence
A tantalizing feature of the spectral $\alpha$-Rényi divergence for $\alpha\in(0,1)$ is that it is robust with respect to outliers in the frequency domain, which has been discussed in [34, 11]. Theorem 1 explains about this robustness as the $\gamma$-divergence induces the robust parameter estimation scheme [7]. Theorem 2 also provides another explanation about the robustness of the spectral Rényi divergence as discussed in this subsection. Let $\mathcal{S}_{\Theta}:=\{S_{\theta}\in\mathcal{S}:\theta\in\Theta\}$ be a parametric spectral model with a parameter space $\Theta\subset\mathbb{R}^{d}$. Consider the minimum spectral $\alpha$-Rényi divergence estimator $\widehat{\theta}_{\alpha}$ given as
@@ 2.2 Robustness of the spectral Renyi divergence
We begin with introducing outliers and contamination in the frequency domain; see Section 8 of [19] for details. Expressing contamination caused by outliers in the frequency domain, we consider a contaminated pilot estimate $\widetilde{I}_{n}^{z,\omega^{*}}$: for $\omega^{*}\in\Omega_{n}$ and $z>0$,
@@
This contamination is motivated by the case when periodic components are not suitably subtracted and are contaminated in observed time series [10, 12, 22], say, $x_{t}^{o}=x_{t}+\sqrt{8\pi z/n}\cos(At)$, $t=1,\ldots,n$ with original zero-mean time series $\{x_{t}\}_{t=1,2,\cdots,n}$. In this case, the periodogram of $x_{t}^{o}$ has approximately the form of
@@
where $\delta_{A}(\omega)=1$ if $\omega=A$ and 0 otherwise.
$=\frac{1}{n}\frac{\alpha S_{\theta}(\omega^{*})}{\alpha S_{\theta}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}\frac{z}{(1-\alpha)z+\alpha S_{\theta}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}$
$\quad-\frac{1}{n}\log(1+\widetilde{I}_{n}^{-1}(\omega^{*})z)+\frac{1}{n}\log(1+(1-\alpha)z\{\alpha S_{\theta}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})\}^{-1})$
$=o_{P}(1)$
3 Optimization paths
@@
This section presents a further robustness property of the minimum spectral Rényi divergence estimator. We show that (1) the optimization for the minimum spectral Rényi divergence estimator is robust against to outliers in the frequency domain, whereas (2) the optimization for the minimum Itakura–Saito divergence estimator is sensitive to them.
3.1 Stable optimization path of the Spectral Rényi divergence minimization